# Pre-training Auto-regressive Robotic Models with 4D Representations

### v1

@@ 1 Introduction
We summarize our main contributions as follows: (i) We introduce a novel robotics pre-training approach that incorporates low-level 4D representations that enhance understanding of the physical world while also learning from unlabeled videos. (ii) Our approach shows that pre-training solely on human video data can lead to better performance than other methods that are pre-trained only on robotic data; (iii) Our method on average surpasses baselines like PerAct (Shridhar et al., 2023) on 12 different tasks in RLBench’s simulated environment, and OpenVLA (Kim et al., 2024) on real tasks with a 7-DoF Kinova Gen3 robot; (iv) Our model also exhibits several advantageous properties, including cross-robot generalization and 3D point track prediction for out-of-domain human and robotic videos.
@@ 4.2 Simulation Evaluation
Method | pick cube up | destack | stack | ||||
yellow | cyan | green | yellow | cyan | yellow on cyan | cyan on yellow | |
ATM | 5.3 3.5 | 6.7 2.7 | 9.3 1.3 | 4.0 2.3 | 9.3 3.5 | 1.3 1.3 | 2.6 1.3 |
OpenVLA | 77.8 6.4 | 45.8 4.2 | 91.7 8.3 | 55.6 2.8 | 51.3 2.6 | 27.8 2.8 | 38.5 4.4 |
Ours | 92.6 3.7 | 100 0.0 | 95.8 4.2 | 94.4 2.7 | 94.9 5.1 | 63.6 5.2 | 59.5 2.4 |
Method | pick toys then place to target | push | Average | ||||
spiderman | penguin | pig | play basketball | push red button | push red the blue | ||
ATM | 5.3 1.3 | 6.7 1.3 | 5.3 3.5 | 24.0 4.6 | 4.0 2.3 | 0.0 0.0 | 6.4 2.2 |
OpenVLA | 2.7 1.3 | 17.3 1.3 | 2.7 2.7 | 49.3 3.5 | 23.1 4.4 | 0.0 0.0 | 37.2 3.4 |
Ours | 90.7 1.3 | 94.7 1.3 | 93.3 1.3 | 92.0 2.3 | 84.6 4.4 | 25.0 4.8 | 83.1 3.0 |
@@ 4.3 Real Robot Evaluation
Baselines. We evaluate our model against two baselines in real-world settings: ATM (Wen et al., 2024) and OpenVLA (Kim et al., 2024). ATM utilizes a hierarchical framework to predict 2D point trajectories, which are then used to condition a policy. In contrast, ARM4R predicts 3D point trajectories, a more intuitive and natural representation for robotic tasks. OpenVLA, a state-of-the-art 7B-parameter VLA model, is pre-trained on the OpenX dataset, while ARM4R is trained on a significantly smaller dataset, with pre-training consisting exclusively of human video data. More details for these implementations are in Section D.1.
Results. Table 2 shows that ARM4R outperforms both baselines across all tasks, achieving an average success rate of 83.1%, compared to OpenVLA’s 37.2% and ATM’s 6.4%. ATM in particular does not perform well in our real setting despite training with a significantly larger number of demonstrations than we use in our fine-tuning. We believe that this significant gap in performance is due to how we track points: ARM4R utilizes 3D coordinates, while ATM relies on 2D. The use of 3D coordinates provides a more natural and accurate representation for robotic tasks, which may contribute to our model’s improved performance.
In contrast to ATM, OpenVLA uses a similar number of fine-tuning episodes to our evaluation setting. However, we believe that our superior performance over OpenVLA can again be attributed to our use of low-level 4D representations, which enable 3D scene understanding.
@@ 4.5 Additional Experiments
We perform additional experiments to evaluate our pre-training effectiveness, and how well the 3D point representations can generalize. More experiments are in Appendix A.

### v2

@@ 1 Introduction
We summarize our main contributions as follows: (i) We introduce a novel robotics pre-training approach that incorporates low-level 4D representations that enhance understanding of the physical world while also learning from unlabeled videos. (ii) Our approach shows that pre-training solely on human video data can lead to better performance than other methods that are pre-trained only on robotic data; (iii) Our method on average surpasses baselines like PerAct (Shridhar et al., 2023) on RLBench, and LLARVA (Niu et al., 2024), OpenVLA (Kim et al., 2024), and -FAST (Pertsch et al., 2025) on real tasks with a 7-DoF Kinova Gen3 robot; (iv) Our model also exhibits several advantageous properties, including cross-robot generalization and 3D point track prediction for out-of-domain human and robotic videos.
@@ 2 Related Work
RVT (Goyal et al., 2023) and RVT-2 (Goyal et al., 2024) are recent transformer-based methods that use a more direct approach for learning 3D information, leveraging ground-truth RGB-D images to reconstruct a scene’s point cloud and then predict keyframes. This point cloud reconstruction relies on a similar principle to our method, with ARM4R instead using pre-training on 3D point tracking to learn scene structure.
@@ 3.1 Preliminaries
4D Representations for Robotic Manipulators. One of the main benefits of using 4D representations is that the 3D point tracks in a robotic setting are described by linear transformations constructed using the robot states. We now present a simple proof of this claim.
Consider a -DoF open-chain manipulator, initially in its reference configuration . Suppose the manipulator is commanded to a new position . Next, we take any arbitrary point on the body of the manipulator lying between joints . Let be described by the transformation in joint ’s frame. Then, ’s new position in the base frame can be described by:
where are the twists associated with each joint. Since this product of 4x4 SE(3) matrices represents a linear transformation, points on the robot body as well as attached rigid bodies evolve through linear transformations in terms of the robot states. Additionally, for overactuated manipulators (7-DoF), the robot learning problem simplifies to tracking arbitrary trajectories, as configuration space limitations are less restrictive. This allows the model to benefit more effectively from the 3D point tracking pre-training task.
@@ 4.2 Simulation Evaluation
Method pick cube up destack stack yellow cyan green yellow cyan yellow on cyan cyan on yellow ATM 5.3 3.5 6.7 2.7 9.3 1.3 4.0 2.3 9.3 3.5 1.3 1.3 2.6 1.3 LLARVA 44.4 6.4 41.6 4.2 54.2 11.0 8.3 4.8 10.3 2.6 5.6 2.8 12.8 2.6 -FAST 63.0 3.7 33.3 8.3 25.0 0.0 5.5 2.8 25.6 2.6 22.2 2.8 25.6 6.8 OpenVLA 77.8 6.4 45.8 4.2 91.7 8.3 55.6 2.8 51.3 2.6 27.8 2.8 38.5 4.4 Ours 92.6 3.7 100 0.0 95.8 4.2 94.4 2.7 94.9 5.1 63.6 5.2 59.5 2.4 Method pick toys then place to target push Average spiderman penguin pig play basketball push red button push red then blue ATM 5.3 1.3 6.7 1.3 5.3 3.5 24.0 4.6 4.0 2.3 0.0 0.0 6.4 2.2 LLARVA 9.3 1.3 9.3 1.3 8.0 2.3 10.7 3.5 20.5 5.1 2.8 2.8 18.3 3.9 -FAST 16.0 2.3 17.3 1.3 9.3 2.7 13.3 3.5 20.5 2.6 0.0 0.0 21.2 3.0 OpenVLA 2.7 1.3 17.3 1.3 2.7 2.7 49.3 3.5 23.1 4.4 0.0 0.0 37.2 3.4 Ours 90.7 1.3 94.7 1.3 93.3 1.3 92.0 2.3 84.6 4.4 25.0 4.8 83.1 3.0
@@
Failure Cases Analysis. We also found some abnormal behaviors when evaluating ARM4R on RLBench, which can be further summarized into two categories. (1) Unnatural rotation: we examined a new task,“put knife” in RLBench and found our model struggles to succeed at this task. Interestingly, this simulated task features an unnatural whole-arm rotation to grasp the knife handle, based on the expert demonstrations created by the simulator’s motion planning. As the movement between two key points is very long, we hypothesize that our model has difficulty learning this rotation. (2) Lack of Precision: we observed that ARM4R struggles with the ‘screw bulb’ task, a procedure requiring precise insertion into the bulb holder. In contrast, our model effectively performs other standard precision tasks, such as ‘open drawer’ which necessitate only standard precision.
@@ 4.3 Real Robot Evaluation
Baselines. We evaluate our model against four baselines in real-world settings: ATM (Wen et al., 2024), LLARVA (Niu et al., 2024), -FAST (Pertsch et al., 2025) and OpenVLA (Kim et al., 2024). ATM utilizes a hierarchical framework to predict 2D point trajectories, which are then used to condition a policy, while LLARVA uses 2D point prediction as an auxiliary task for learning action prediction. In contrast, ARM4R predicts 3D point trajectories, a more intuitive and natural representation for robotic tasks. -FAST and OpenVLA are state-of-the-art VLA models pre-trained on the OpenX dataset, while ARM4R is trained on a significantly smaller dataset, with pre-training consisting exclusively of human video data. More details for these implementations are in Section D.1.
Results. Table 2 shows that ARM4R outperforms all baselines across all tasks, achieving an average success rate of 83.1%, compared to OpenVLA’s 37.2% and ATM’s 6.4%. ATM in particular does not perform well in our real setting despite training with a significantly larger number of demonstrations than we use in our fine-tuning. We believe that this significant gap in performance is due to how we track points: ARM4R utilizes 3D coordinates, while ATM relies on 2D. The use of 3D coordinates provides a more natural and accurate representation for robotic tasks, which may contribute to our model’s improved performance.
In contrast to ATM, OpenVLA and -FAST use a similar number of fine-tuning episodes to our evaluation setting. However, we believe that our superior performance over these baselines can again be attributed to our use of low-level 4D representations, which enable 3D scene understanding.
@@ 4.5 Additional Experiments
We conduct additional experiments to assess the benefits of pre-training, generalization of 3D point representations, and robustness to noise. More experiments are in Appendix A.
@@
Robustness Analysis. We conduct an experiment to test how ARM4R handles dynamic changes in its environment. We test the model on three cube tasks (Pick, Stack, and Destack) in our real Kinova setting, manually moving the cube during rollout when the arm reaches 1/3 and 2/3 of its descent. The results are presented in Table 5, and show consistent performance despite movement of the target object. This highlights the model’s robustness to environmental and object-level changes, which we believe is partly due to the fact that it predicts each step in a trajectory and thus interacts with the environment continuously.
Disturbance Time | Pick | Stack | Destack |
No Disturbance | 96.02.3 | 61.31.3 | 94.71.3 |
At 1/3 Descent | 94.73.5 | 60.00.0 | 93.31.3 |
At 2/3 Descent | 92.02.3 | 57.31.3 | 90.71.3 |
To further assess ARM4R’ robustness in dealing with factors like camera noise, occlusions, and sensor drift, we perform additional evaluations on the three cube tasks under the following conditions: (1) Dim lighting: ambient light is reduced to 50%, (2) Background distractors: dynamic background changes, such as people walking by or movement of background curtains, and (3) Tabletop distractors: two random objects are placed near the target. The results are presented in Table 6. Overall, ARM4R demonstrates robustness to lighting and background changes. This is likely due to the use of attention pooling, which guides the model to focus on the target region rather than background features. This hypothesis is supported by the observed performance drop when tabletop distractors are introduced.
Noise Source Pick Stack Destack Standard 96.02.3 61.31.3 94.71.3 Dim Light 86.71.3 52.02.3 81.32.7 Background Distractor 94.71.3 57.31.3 90.71.3 Table-top Distractors 81.31.3 48.02.3 74.71.3
@@ D.1 Baselines in Real Experiments
LLARVA. We reproduce LLARVA (Niu et al., 2024) as a baseline for our real-world Kinova experiment setup, following the provided code and instructions444https://github.com/Dantong88/LLARVA. We fine-tune their released pre-trained checkpoint on our Kinova demonstrations for single-task policies (5 tasks, each with 200 episodes per variation). We use their default learning rate and batch size, and train for 4 epochs for each task.
-FAST. We fine-tune -FAST as a baseline for our Kinova real-world experiment setup, following the provided code and instructions555https://github.com/Physical-Intelligence/openpi. We use their released base autoregressive -FAST model for fine-tuning a single-task policy (200 demonstrations per variation). We use joint position control instead of joint velocity control, fine-tuning the pre-trained model to predict the absolute 7-DoF joint pose plus 1-DoF gripper status. We follow the default learning rate used in their implementation.